OpenStudy (adamaero):

Easy MEDAL - Bob gets a CD for a year, and it returns $5. Sally invests in stocks for a year, and gets a return of $500. What is the percent difference?

OpenStudy (adamaero):

500/5=100 100*100% = 10,000% ?????????????????????????

OpenStudy (wolf1728):

It seems impossible to calculate this. You are not told the amount either one invests. You are not told the interest rate.

OpenStudy (adamaero):

$10,000. I said the returns for each.

OpenStudy (adamaero):

I don't see how the money before the returns is connected to the percent increases BETWEEN the returns.

OpenStudy (wolf1728):

For simple interest: Total = (1+rate) * 10,000 Bob (10,005/10,000) = 1+rate 1.0005 = 1+rate rate = = .0005 = .05%

OpenStudy (adamaero):

Maybe this is what's confusing you: CD = certificate of deposit.

OpenStudy (wolf1728):

Well, if I'm confused, guess I'll just leave then. I do know about compound interest but you should phrase your questions a lot better.

OpenStudy (adamaero):

This is just about sheer differences expressed as a percent (nothing to do with compounding)